OpenStudy (anonymous):

solve by substitution method 3x + 4y = 7 -9x + y = 31 the solution is ---- type an ordered pair or there is no solution

OpenStudy (amistre64):

(-3,4) is what my ti83 practice tells me :)

OpenStudy (amistre64):

\[\left[\begin{array}c 3&4&|&7\\-9&1&|&31\end{array}\right]\] \[\left[\begin{array}c 1&0&|&-3\\0&1&|&4\end{array}\right]\]

OpenStudy (anonymous):

so my answer is -3 and 4 which is the ordered pair i messed up along the second half thanks

OpenStudy (amistre64):

yep; the substitution method is alot of effort; but when you keep track of it all, its an intuitive method

OpenStudy (anonymous):

thanks again i wanna be your fan